Solve using substitution
3x+2y=6
x=34-6y

Found 2 solutions by mananth, mljoo:
Answer by mananth(16949)

3x+2y=6
x=34-6y
..
plug the value of x in the first equation
3(34-6y)+2y=6
102-18y+2y=6
-16y=-102+6
-16y=-96
y=-96/-16
y=6
..
plug the value of y in the same equation
3x+2*6=6
3x+12=6
3x=-12+6
3x=-6
x=-4

Answer by mljoo(4)

3x+2y=6
x=34-6
3(34-6)+2y=6
102-18+2y=6 transpose all the like terms
2y= 6-102+18
2y= 114
2 = 2
cancel the 2
y= 57
3x+2(57)= 6
3x+114=6 transpose the like terms
3x= 6-114
3x= 108
3 = 3
cancell the three
x= 36